Hi,my name is Maria.I have tried to follow the book and still do not understand how they come up with the answer.Question is $50 deposited 100years ago has the value of $1600.today. What was the interest rate? I tried to follow the formula of (I+R)^n*amount.I put in different percentages.ex: (1=.06)^100*50=16965.Am I anywhere close to an answer? Please help!!
Thanks Maria

Answer by nerdybill(7384)

Question is $50 deposited 100years ago has the value of $1600.today. What was the interest rate? I tried to follow the formula of (I+R)^n*amount.
"Simple compound" interest formula:
A = P(1 + r)^n
where
A is amount ($) after n years
P is principal (amount started with)
r is interest rate
n is time (years)
.
Plugging in what was given:
1600 = 50(1 + r)^100
now we solve for r:
1600 = 50(1 + r)^100
1600/50 = (1 + r)^100
32 = (1 + r)^100
32^(1/100) = 1 + r
1.0353 - 1 = r
.0353 = r
or
3.5% = r